# Jiakang Yuan1,2, Xiangchao Yan2, Botian Shi2, Tao Chen1,🖂, Wanli Ouyang2 Bo Zhang2,‡,🖂, Lei Bai2,🖂, Yu Qiao2, Bowen Zhou2

### v2

@@ 3.2 Experimental Verification Process
“Jump in and get your feet wet.” In the research cycle, experimental verification is crucial as it serves as the most effective way to confirm the validity of a proposed idea. Most of the existing automatic scientific discovery works [20, 36] directly evaluate the novelty of AI-generated ideas by LLMs or humans. However, although novel ideas can be obtained in this way, their effectiveness cannot be guaranteed due to the lack of closed-loop experimental verification, which is inconsistent with our scientific research goals. In contrast, Dolphin screens out truly effective ideas through an experimental verification process.
@@
Based on our observation, we design an exception-traceback-guided debugging process as shown in Fig. 3, aiming at assisting the LLMs in comprehending code logic with local code structure. Specifically, to generate the code structure related to the code errors, Dolphin first extracts information in exception traceback, including function name, line, and code, since traceback contains the nested information between functions. Note that Dolphin only focuses on custom codes, excluding library function calls. Then, Dolphin prompts the LLM to generate the code structure under the guidance of extracted exception traceback information. After that, the LLM analyzes the exception traceback and local code structure to make necessary modifications, enabling automatic code execution after these adjustments. The debugging process will be repeated until either successful execution is achieved or the predetermined maximum number of debugging times is reached. Subsequently, all successfully implemented ideas will undergo comprehensive analyses in the next phase.
@@ 3.3 Results Feedback Process
“Experience is the best teacher.” Human researchers often analyze experimental results to further propose new ideas or improve existing ideas, since insights from previous experiments can be leveraged to effectively enhance the quality of subsequent idea generation. However, recent works either implement feedback mechanisms within the isolated idea generation process [20] or lack feedback mechanisms entirely [23]. To address this limitation, Dolphin analyzes experimental results and incorporates the findings into the subsequent round of ideas generation process.
Dolphin first divides the experimental results into three categories (i.e., improvement, maintenance, and decline) compared to the performance of the reference codes. Our goal is to discourage the development of ideas that have previously led to stagnant or declining performance, while actively promoting the creation of innovative concepts or iterations based on past ideas that enhance the model performance. In detail, Dolphin incorporate the embeddings of summaries from ideas that maintain or improve the performance into $\mathbf{B}$ defined in Sec. 3.1. In this way, ideas will be filtered out if they are similar to previous ideas that cannot improve the performance and avoid redundant verification of the ineffective ideas. Besides, Dolphin incorporates performance-enhancing ideas into the idea generation prompt for the next loop. Please refer to the supplementary material for detailed prompts.
@@ 4 Experiment
Tasks CIFAR-100 ModelNet40 SST-2
Top-1 Acc. (%) OA (%) mAcc. (%) Acc. (%)
Baseline 79.5 (WRN-28-10 [49]) 89.2 (PointNet [28]) 86.2 (PointNet [28]) -
Baseline 81.2 (WRN-28-10) 91.0 (PointNet [28]) 87.6 (PointNet [28]) 91.0 (BERT-base [8])
Avg. Improvement 81.8 (+0.6) 92.0 (+1.0) 88.7 (+1.1) 91.8 (+0.8)
Max Improvement 82.0 (+0.8) 93.9 (+2.9) 91.1 (+3.5) 92.5 (+1.5)
Human designed 82.2 (ResNeXt [22]) 93.8 (GPSFormer [40]) 91.8 (GPSFormer [40]) 93.1 (BERT-large [8])
Number ideas 6 / 40 5 / 40 5 / 40 6 / 40
Table 1: Experimental verifications on 2D image classification, 3D point classification, and sentiment classification tasks. Number ideas refers to the number of ideas that can achieve performance gains. $\dagger$ denotes the results of our implementation. Avg. Improvement and Max Improvement represent the average and maximum improvement of all ideas that can improve the baseline performance.
@@ 4.1 Experimental Setups
Tasks and Datasets. We conduct auto-research on three topics including image classification, 3D classification, and sentiment classification. For 2D classification task, we evaluate our method on CIFAR-100 [19], which is widely-used in computer vision. For 3D classification task, we use ModelNet40 [45] which is a 3D CAD dataset and consist of 40 categories. For the sentiment classification task, we use Stanford Sentiment Treebank (SST-2) dataset [35]. More details can be found in supplementary materials.
Implementation Details. For the ideas generation process, we use gpt-4o-2024-08-06 [25] as our LLM agent. The total number of retrieved papers is set to 50 and only papers with scores higher than 8 will be treated as references for the ideas generation process. We generate 20 ideas in each loop and the threshold of the independence filtering is set to 0.8. We use sentence-transformer/all-roberta-large-v1 [33] to extract the summary embedding of each idea. For the experimental verification process, we use deepseek-v2.5 [53] deployed by ollama [39] as our code agent. The maximum number of debugging attempts is set to 5 for experimental efficiency. Following AI-Scientist [23], we use self-reflection to first eliminate some syntax errors before executing the program and use aider as the framework to call LLM agents. The same hyper-parameters and models employed in the ideas generation process are utilized in the results feedback.
@@ 4.2 Main Results
We evaluate Dolphin’s capabilities across various tasks covering images, point clouds, and language modalities. In this section, we conduct experiments on each task for two loops (i.e., 40 ideas). For more detailed analyses of the closed-loop process, please refer to Sec. 4.3. Detailed implementations for the different tasks can be found in our supplementary material.
Results on 2D Image Classification. We first conduct experiments on the image classification task, using WRN-28-10 [49] as our baseline model. This model is trained on CIFAR-100 [19] for 200 epochs, and we report the Top-1 accuracy. As shown in Tab. 1, the average improvement and max improvement are 0.6% and 0.8%, respectively. Notably, the idea generated and performed automatically by Dolphin can achieve comparable performance to hand-crafted methods such as ResNeXt [22] (e.g., 82.0% compared to 82.2%). It should be noted that Transformer-based methods [9] such as ViT are not included in our comparison, due to their heavy dependence on pre-training on large-scale datasets, which, at this stage, requires significant resource investment to validate their effectiveness.
@@
Results on 3D Point Classification. Further, we conduct experiments on 3D classification task using PointNet [28] as our baseline. We train the model on ModelNet40 [45] for 200 epochs, and report both overall accuracy and mean accuracy. As illustrated in Tab. 1, a total of 5 ideas achieve performance gains in two loops, with an average improvement of 1.0% OA, which shows that Dolphin can generate effective ideas and verify the idea by performing experiments. Besides, the maximum improvement can achieve 93.9% overall accuracy, which largely improves the performance compared with the human-designed baseline (i.e., 91.0% achieved by PointNet [28]). More excitingly, such results achieved by auto-research achieve comparable performance to the current state-of-the-art method (i.e., 93.8% achieved by GPSFormer [40]). This method is carefully designed by human researchers based on Transformer architecture. We would like to emphasize that for a fair comparison, we compare the 3D methods without pre-training and the voting mechanism.
@@
Results on Sentiment Classification. To verify the effectiveness of Dolphin on different modalities, we also perform experiments on the NLP task (i.e., sentiment classification). We conduct experiments on SST-2 [35] and report the classification accuracy. We fine-tune the pre-trained BERT-base model [8] as our baseline. It can be seen that Dolphin can also generate and perform effective ideas (e.g., 1.5% improvement on SST-2) on NLP task, reducing the performance gap between BERT-base (i.e., 91.0%) and BERT-large (i.e., 93.1%).
4.3 Further Analyses
@@
Analysis on Ideas Generation Process. To evaluate the effectiveness of the ideas generation process. We conduct further ablation studies by comparing naive generation, generation with retrieved papers, and our proposed method. Naive generation refers to the direct use of LLMs to generate ideas based on the seed idea and reference codes, similar to the approach used by AI-Scientist [23]. Generation with retrieved papers involves directly searching papers based on the topic and filtering them by their relevance to the input topic. As shown in Tab. 2, the naive generation yields the poorest results, with more than half of the ideas being judged as not novel. Furthermore, the quality of generated ideas is significantly improved when using naive retrieved papers, as this approach more closely aligns with the way human researchers generate ideas. However, as mentioned in Sec. 3.1, this approach tends to retrieve irrelevant papers and will mislead LLMs. As indicated in the first line of Tab. 3, some papers primarily focus on 3D detection or point cloud completion, where the design approach of the model is entirely different from that of the 3D point classification. This phenomenon can be well handled by the designed paper ranking process. As illustrated in Tab. 2 and Tab. 3, the number of novel ideas significantly increased from 8/20 to 19/20, while the proportion of papers related to irrelevant topics substantially decreased. This improvement can be attributed to the availability of more relevant reference papers during the ideas generation process. It is worth noting that the higher occurrence of the keyword “segmentation” is due to that, many studies concurrently perform both point classification and segmentation tasks. In addition, the average cost per idea is shown in Tab. 2. It can be seen that the cost of each idea is very small. Besides, the cost of generating each idea is relatively higher when retrieving papers. This is mainly due to the retrieval process and the longer prompt required for idea generation (i.e., both the title and abstract are fed into the LLMs as references).
@@
PointNet [28] 86.2 89.2
PointNet++ [29] - 91.9
DGCNN [43] 90.2 92.9
PointNeXt [30] 90.8 93.2
OctFormer [41] - 92.7
GPSFormer [40] 91.8 93.8
@@
Table 6: Classification on ModelNet40 [45]. The results are obtained from 1024 points without voting.
@@ 4.4 Case Studies
Diff. DGCNN PointNet-CSR (Completed by Dolphin)
@@
Table 7: The differences between DGCNN [43] proposed by human and PointNet-CSR proposed using Dolphin.
We illustrate our approach with an example drawn from 3D point classification task as shown in Fig. 4. It can be seen that Dolphin can generate codes corresponding to the idea and only in this way, the generated ideas can be effectively verified. Tab. 6 presents the comparison between AI-generated approach (i.e., PointNet-CSR obtained by Dolphin) and previous human-designed methods. Idea automatically generated and performed by Dolphin can outperform most human-designed methods and achieve comparable performance to current SOTA [40]. Furthermore, as shown in Tab. 7, we carefully investigate the existing works on 3D classification task up to the submission date, identifying the work most relevant to PointNet-CSR (AI-generated 3D work), as illustrated in Fig. 4. The detailed comparison of the idea, implementation, and result can be found in Tab. 7, showing that PointNet-CSR can achieve better and faster performance through a more concise architecture. Please refer to our supplementary material for more comparisons between human-designed works and Dolphin-generated works.
@@ 5 Conclusion and Future Outlook
Appendix
Outlines
In the Appendix section, we provide additional details and qualitative results from the following aspects:
  • Sec. 6: More details of Dolphin including prompts, qualitative results of each step, and so on:
    • Sec. 6.1: Ideas generation process.
    • Sec. 6.2: Experimental verification process.
  • Sec. 7: More details and results of performing experiments using Dolphin:
    • Sec. 7.1 The selected datasets and implementation on different tasks:
        Sec. 7.1.1: Image classification task.
        Sec. 7.1.2: 3D classification task.
        Sec. 7.1.3: Sentiment classification task.
    • Sec. 7.2: Code implementation differences between human-designed and Dolphin-designed.
  • Sec. 8: More analyses and further works:
    • Sec. 8.1: The extensibility of the research task related to idea generation.
    • Sec. 8.2: Analyses on future works.
6 More details of Dolphin
In this section, we provide more details about Dolphin including prompts, qualitative results of some processes, algorithms of some processes, and so on. In the following section, we will give more details of the ideas generation process, experimental verification process, and results feedback process, respectively.
6.1 Ideas Generation Process
We provide prompts used in paper retrieval, paper ranking, and idea generation process in Fig. 5. We partially refer to the prompt design of previous works [23, 34]. As depicted in our manuscript, the quality of retrieved papers is important to idea generation. Here, we give an example to further illustrate the impact. Given the topic ‘3D classification’, naive retrieval will result in lots of papers related to 3D object detection. As a result, we have identified several ideas that are more closely related to detection (e.g., region proposal PointNet), which are significantly influenced by the detection task.

### v3

@@ 3.2 Experimental Verification Process
“Jump in and get your feet wet.” Experimental verification is crucial in the research cycle as it serves as the most effective way to confirm the effectiveness of proposed ideas. Most existing works Li et al. (2024); Su et al. (2024) assess the novelty of AI-generated ideas via LLMs or humans. However, without closed-loop experimental verification, the effectiveness of these ideas remains uncertain. In contrast, Dolphin screens out truly effective ideas through an experimental verification process.
@@
Based on our observation, we design an exception-traceback-guided debugging process as shown in Fig. 3, aiming at assisting the LLMs in comprehending code logic with local code structure. Specifically, to generate the code structure related to the code errors, Dolphin first extracts information in exception traceback, including function name, line, and code, since traceback contains the nested information between functions. Note that Dolphin only focuses on custom codes, excluding library function calls. Then, Dolphin prompts the LLM to generate the code structure under the guidance of extracted exception traceback information. After that, the LLM analyzes the exception traceback and local code structure to make necessary modifications, enabling automatic code execution after these adjustments. The debugging process will be repeated until either successful execution is achieved or the predetermined maximum number of debugging times is reached. Finally, all successfully implemented ideas will undergo comprehensive analyses in the next phase.
@@ 3.3 Results Feedback Process
“Experience is the best teacher.” Human researchers often analyze experimental results to further propose new ideas or improve existing ideas, since insights from previous experiments can be leveraged to effectively enhance the quality of subsequent idea generation. However, recent works either implement feedback mechanisms within the isolated idea generation process Li et al. (2024) or lack feedback mechanisms entirely Lu et al. (2024). To address this limitation, Dolphin analyzes experimental results and incorporates the findings into the subsequent round of ideas generation process.
Dolphin first divides the experimental results into three categories (i.e., improvement, maintenance, and decline) compared to the performance of the reference codes. Our goal is to discourage the development of ideas that have previously led to stagnant or declining performance, while actively promoting the creation of innovative concepts or iterations based on past ideas that enhance the model performance. In detail, Dolphin incorporate the embeddings of summaries from ideas that maintain or improve the performance into $\mathbf{B}$ defined in Sec. 3.1. In this way, ideas will be filtered out if they are similar to previous ideas that cannot improve the performance and avoid redundant verification of the ineffective ideas. Besides, Dolphin incorporates performance-enhancing ideas into the idea generation prompt for the next loop. Detailed prompts are shown in Appendix A.
Tasks ModelNet40 CIFAR-100 SST-2
OA (%) mAcc. (%) Top-1 Acc. (%) Acc. (%)
Baseline 89.2 (PointNet) 86.2 (PointNet) 79.5 (WRN-28-10) -
Baseline 91.0 (PointNet) 87.6 (PointNet) 81.2 (WRN-28-10) 91.0 (BERT-base)
Avg. Improvement 92.0 (+1.0) 88.7 (+1.1) 81.8 (+0.6) 91.8 (+0.8)
Max Improvement 93.9 (+2.9) 91.1 (+3.5) 82.0 (+0.8) 92.5 (+1.5)
Human designed 93.8 (GPSFormer) 91.8 (GPSFormer) 82.2 (ResNeXt) 93.1 (BERT-large)
Number ideas 5 / 40 5 / 40 6 / 40 6 / 40
Table 1: Experimental verifications on 3D point classification, 2D image classification, and sentiment classification tasks. Number ideas refers to the number of ideas that can achieve performance gains. $\dagger$ denotes the results of our implementation. Avg. Improvement and Max Improvement represent the average and maximum improvement of all ideas that can improve the baseline performance.
@@ 4.1 Experimental Setups
Tasks and Datasets. We conduct auto-research on three topics including image classification, 3D classification, and sentiment classification. For 2D classification task, we evaluate our method on CIFAR-100 Krizhevsky et al. (2009), which is widely-used in computer vision. For 3D classification task, we use ModelNet40 Wu et al. (2015) which is a 3D CAD dataset and consist of 40 categories. For the sentiment classification task, we use Stanford Sentiment Treebank (SST-2) dataset Socher et al. (2013). We also conduct experiments on MLE-bench Chan et al. (2024), which is designed to evaluate the capability of models in handling real-world ML tasks on Kaggle competitions. Further details of datasets can be found in Appendix B.1.
Implementation Details. For the ideas generation process, we use gpt-4o-2024-08-06 OpenAI (2024) as our LLM agent. The total number of retrieved papers is set to 50 and only papers with scores higher than 8 will be treated as references for the ideas generation process. We generate 20 ideas in each loop and the threshold of the independence filtering is set to 0.8. We use sentence-transformer/all-roberta-large-v1 Reimers and Gurevych (2020) to extract the summary embedding of each idea. For the experimental verification process, we use deepseek-v2.5 Zhu et al. (2024) deployed by ollama Team (2023b) as our code agent. The maximum number of debugging attempts is set to 5 for experimental efficiency. Following AI-Scientist Lu et al. (2024), we use self-reflection to first eliminate some syntax errors before executing the program and use aider as the framework to call LLM agents. The same hyper-parameters and models employed in the ideas generation process are utilized in the results feedback.
@@ 4.2 Main Results
We evaluate Dolphin’s capabilities across various tasks covering point clouds, images, and language modalities. In this section, we conduct experiments on each task for two loops (i.e., 40 ideas).
@@
Results on 3D Point Classification. We conduct experiments on 3D classification task using PointNet Qi et al. (2017a) as our baseline model. As illustrated in Tab. 1, a total of 5 ideas achieve performance gains in two loops, with an average improvement of 1.0% OA, which shows that Dolphin can generate and verify effective ideas. Besides, the maximum improvement can achieve 93.9% OA, which largely improves the performance compared with the human-designed baseline (i.e., 91.0% achieved by PointNet). More excitingly, such results achieved by auto-research obtain comparable performance to the current SoTA method (i.e., 93.8% achieved by GPSFormer Wang et al. (2024a)). This method is carefully designed by human researchers based on Transformer architecture. We would like to emphasize that for a fair comparison, we compare the 3D methods without pre-training and the voting mechanism.
Results on 2D Image Classification. Further, we conduct experiments on the image classification task, using WRN-28-10 (Zagoruyko, 2016) as our baseline model. As shown in Tab. 1, the average improvement and max improvement are 0.6% and 0.8%, respectively. Notably, the idea generated and performed automatically by Dolphin can obtain comparable performance to hand-crafted methods such as ResNeXt (Liu et al., 2022) (e.g., 82.0% compared to 82.2%). It should be noted that Transformer-based methods such as ViT (Dosovitskiy et al., 2020) are not included in our comparison, due to their heavy dependence on large-scale pre-training, which, at this stage, requires significant resources to validate their effectiveness.
@@
Results on Sentiment Classification. To verify the effectiveness of Dolphin on different modalities, we also perform experiments on sentiment classification task. We conduct experiments on SST-2 Socher et al. (2013) and report the classification accuracy. In Tab. 1, we fine-tune the pre-trained BERT-base Devlin et al. (2019) as our baseline model. It can be seen that Dolphin also generates and performs effective ideas (e.g., 1.5% improvement on SST-2) on sentiment classification, reducing the performance gap between BERT-base (i.e., 92.5%) and BERT-large (i.e., 93.1%).
Results of tasks in MLE-bench. We further conduct experiments on tasks from MLE-bench Chan et al. (2024) as shown in Tab. 2. First, Dolphin can flexibly integrate with code generation frameworks such as AIDE Dominik Schmidt (2024) and Agent Laboratory Schmidgall et al. (2025). For example, based on the code template generated by AIDE, Dolphin further improves the performance on tasks such as insult detection (i.e., 81.0% $\rightarrow$ 84.7%) and achieves gold medal level performance on Kaggle. Additionally, Dolphin performs version updates for technologies or code. For example, based on previous code from Kaggle, Dolphin can achieve further accuracy breakthroughs (e.g., 95.3 $\rightarrow$ 96.2 on tabular playground).
Tasks Domain Code src. Prev. score Score
[…]
Table 2: Results of tasks in MLE-bench. Code src. denotes the source of the code template (AIDE: auto-generate by AIDE, Kaggle: Kaggle public code).
@@
4.3 Further Analyses
Analysis on Ideas Generation Process. To evaluate the effectiveness of the ideas generation process, we conduct further ablation studies by comparing naive generation, generation with retrieved papers, and our proposed method. Naive generation refers to the direct use of LLMs to generate ideas based on the seed idea and reference codes, similar to the approach used by AI-Scientist Lu et al. (2024). Generation with retrieved papers involves directly searching papers based on the topic and filtering them by their relevance to the input topic. As shown in Tab. 3, the naive generation yields the poorest results, with more than half of the ideas being judged as not novel. Furthermore, the quality of generated ideas is significantly improved when using naive retrieved papers, as this approach more closely aligns with the way human researchers generate ideas. However, as mentioned in Sec. 3.1, this approach tends to retrieve irrelevant papers and will mislead LLMs. As indicated in the first line of Tab. 4, some papers primarily focus on 3D detection or point cloud completion, where the design approach of the model is entirely different from that of point classification. This phenomenon can be well handled by the designed paper ranking process. As illustrated in Tab. 3 and Tab. 4, the number of novel ideas significantly increased from 8/20 to 19/20, while the proportion of papers related to irrelevant topics substantially decreased. This improvement can be attributed to the availability of more relevant reference papers during the ideas generation process. Note that the higher occurrence of the keyword “segmentation” is due to that, many studies concurrently perform both point classification and segmentation tasks. In addition, the average cost per idea is shown in Tab. 3. It can be seen that the cost of each idea is very small. Besides, the cost of generating each idea is relatively higher when retrieving papers. This is mainly due to the retrieval process and the longer prompt required for idea generation (i.e., both the title and abstract are fed into the LLMs as references).
@@ 4.3 Further Analyses
PointNet Qi et al. (2017a) 86.2 89.2
PointNet++ Qi et al. (2017b) - 91.9
DGCNN Wang et al. (2019) 90.2 92.9
PointNeXt Qian et al. (2022) 90.8 93.2
OctFormer Wang (2023) - 92.7
GPSFormer Wang et al. (2024a) 91.8 93.8
@@
Table 7: Accuracy on ModelNet40. The results are obtained from 1024 points without voting.
@@ 4.4 Case Studies
Diff. DGCNN (Human) PointNet-CSR (Dolphin)
@@
Table 8: The differences between DGCNN Wang et al. (2019) proposed by human and PointNet-CSR proposed using Dolphin.
We illustrate our approach with an example drawn from 3D point classification task as shown in Fig. 9. It can be seen that Dolphin can generate codes corresponding to the idea and only in this way, the generated ideas can be effectively verified. Tab. 7 presents the comparison between AI-generated approach (i.e., PointNet-CSR obtained by Dolphin) and previous human-designed methods. Idea automatically generated and performed by Dolphin can outperform most human-designed methods and achieve comparable performance to current SoTA Wang et al. (2024a). Furthermore, as shown in Tab. 8, we carefully investigate the existing works on 3D classification task up to the submission date, identifying the work most relevant to PointNet-CSR (AI-generated 3D work), as illustrated in Fig. 9. The detailed comparison of the idea, implementation, and result can be found in Tab. 8, showing that PointNet-CSR can achieve better and faster performance through a more concise architecture. Please refer to Fig. 9 and Appendix B.2 for more comparisons between human-designed works and Dolphin-generated works.
@@ 5 Conclusion and Future Outlook
Limitations
Although Dolphin can generate and implement effective ideas in a loop, it still has inherent limitations. First, in the idea generation stage, LLMs retain historical knowledge obtained from training data, which may cause knowledge leakage when generating ideas. Besides, only the abstract and title are used in the idea generation process, which may result in LLM not being able to understand the technical details in the article and the logic between the articles. However, the feedback-driven idea generation process in Dolphin can be flexibly combined with methods that address these concerns. In addition, the code capabilities of LLMs are not sufficient to understand complex codes such as project-level codes, resulting in Dolphin being unable to verify complex tasks currently.
Ethics Statement
This work focuses on the development of a closed-loop framework for auto-research, aiming to accelerate the research cycle. The proposed method relies on publicly available datasets, research papers, and models, ensuring compliance with copyright and intellectual property laws. While the framework is intended to aid human research, we encourage users to critically evaluate the generated results to ensure that they adhere to ethical research practices and mitigate any potential limitations, such as bias or incomplete ideas.
        Appendix. B.1.1: Image classification task.
        Appendix B.1.2: 3D classification task.
        Appendix B.1.3: Sentiment classification task.
        Appendix B.1.4: MLE-bench.
    • Appendix B.2: Code implementation differences between human-designed and Dolphin-designed.
  • Appendix C: More analyses and further works:
    • Appendix C.1: The extensibility of the research task related to idea generation.
    • Appendix C.2: Analyses on future works.
Appendix A Further details of Dolphin
In this section, we provide more details about Dolphin including prompts, qualitative results of some processes, and algorithms of some processes. In the following section, we will give more details of the ideas generation process, experimental verification process, and results feedback process, respectively.
A.1 Ideas Generation Process
We provide prompts used in paper retrieval, paper ranking, and idea generation process in Fig. 5. We partially refer to the prompt design of previous works Lu et al. (2024); Si et al. (2024). As depicted in our manuscript, the quality of retrieved papers is important to idea generation. Here, we give an example to further illustrate the impact. Given the topic ‘3D classification’, naive retrieval will result in lots of papers related to 3D object detection. As a result, we have identified several ideas that are more closely related to detection (e.g., region proposal PointNet), which are significantly influenced by the detection task.